OpenStudy (anonymous):

An object falls freely from rest near the surface of Earth, What is the speed of the object after having fallen a distance at 4,90 meters

OpenStudy (anonymous):

Whenever working out equations like this you are taking of model of a real life situation so i'm assuming that you are ignoring air resistance all buoyant forces and terminal velocity as a result of ignoring the previous two? In which case it is again suvat, noting down your information: u=0 a=9.8 (gravity, positive as it is in the same direction as the particle) s=490 v=? using the equation v^2=u^2 + 2as you get v^2 =2*9.8*490 = 9,604 (remember this is v^2) so \[v=\sqrt{9604}=98\] ms^-1 Hope that helped!! You should learn your suvat equations!

OpenStudy (anonymous):

lol i know i have no idea what a suvat equation is though we never learned it yet these questions were on a test